OpenStudy (anonymous):

Question:√3*√21 Answers: √63 √24 None of the above 3√7

OpenStudy (angela210793):

=3sqr7

OpenStudy (angela210793):

yea Raheen but 63=9*7 right?

OpenStudy (anonymous):

3√7

OpenStudy (anonymous):

yes, it's simplified as 3√7 but what's wrong with √63

OpenStudy (angela210793):

nothing's wrong with it but 3sqr7 is better...my teacher wouldn't give me a single point if i gave the answer sqr63..so i simplify them as much as possible :):) High school trauma :P:P

OpenStudy (anonymous):

\[\sqrt{3}\times \sqrt{21}=\sqrt{63}=\sqrt{3^{2}\times7}=3\sqrt{7}\]

OpenStudy (anonymous):

angela you are right

OpenStudy (angela210793):

:)

OpenStudy (anonymous):

"simplest radical form" does not allow \[\sqrt{63}=\sqrt{3^2\times 7}\] because you can't have a square inside a square root to be in that form. that is why you get no points, even though the numbers are identical